# Return of EM: Entity-driven Answer Set Expansion for QA Evaluation

### v1

@@ 4.2 Evaluation Methods
Prior works have also explored answer expansion in different contexts. They use either Freebase (Si et al., 2021), for NQ and TQ datasets or Wikipedia for the TQ dataset (Joshi et al., 2017) [^6] We utilized the Wikipedia-based expansion version of TriviaQA as released by the original authors. However, there is no publicly available code to adapt this method to the NQ.. Our InstructGPT-based expansions employ variants like Inst-zero (only-instruction), Inst-random (randomly selected few-shot), and Inst-entity (entity type-specific few-shot as detailed in Section 3). Based on the expanded answer set, soft EM is used to assess the model predictions. The accuracy of these methods is determined by comparison against human annotations.
@@ 4.3 Results & Analysis
Figure 3 details the specific accuracies of different expansion methods by entity type. Numeric entities and the N/A category, which typically exhibit diverse surface forms, show lower accuracy with the original answer set. Our method significantly improves accuracy for these categories by effectively handling their diversity.
Among the answer set expansion baselines, Wikipedia-based expansion shows competitive performance in non-numeric entity types, leading to high performance in TQ since the majority of data (81.4%) is non-numeric type. However, in numeric entity types, its performance is less effective. This can be attributed to the fact that Wikipedia entities are mostly related to non-numeric types, such as PERSON, LOC, and ORG.
@@
Human judgement: Incorrect
Insteval judgement: Correct
@@
Human judgement: Correct
Insteval judgement: Incorrect
Table 3: Error instances from Insteval, direct evaluation using InstructGPT. The examples are taken from NQ and ten samples from each of the five QA models.

### v2

@@ 4.2 Evaluation Methods
Prior works have also explored answer expansion in different contexts. They use either Freebase (Si et al., 2021), for NQ and TQ datasets or Wikipedia for the TQ dataset (Joshi et al., 2017) [^6] We utilized the Wikipedia-based expansion version of TriviaQA as released by the original authors. However, there is no publicly available code to adapt this method to the NQ.. Our InstructGPT-based expansions employ variants like Inst-zero (only-instruction), Inst-random (randomly selected few-shot), and Inst-entity (entity type-specific few-shot as detailed in Section 3). Based on the expanded answer set, soft EM is used to assess the model predictions.
The accuracy of these methods is determined by comparison against human annotations.
@@ 4.3 Results & Analysis
Figure 3 details the specific accuracies of different expansion methods by entity type. Numeric entities and the N/A category, which typically exhibit diverse surface forms, show lower accuracy with the original answer set. Our method significantly improves accuracy for these categories by effectively handling their diversity. Among the answer set expansion baselines, Wikipedia-based expansion shows competitive performance in non-numeric entity types, leading to high performance in TQ since the majority of data (81.4%) is non-numeric type. However, in numeric entity types, its performance is less effective. This can be attributed to the fact that Wikipedia entities are mostly related to non-numeric types, such as PERSON, LOC, and ORG.
The effectiveness of our method can vary depending on the entity’s popularity (Mallen et al., 2022) or rarity (Kandpal et al., 2023), as it relies on InstructGPT’s background knowledge for each entity. Following  Kandpal et al. (2023), we also present the effectiveness of our method compared to Soft EM with the original answer set, segmented by the rarity of each entity. As shown in Figure 4, our method consistently maintains its effectiveness across varying levels of entity rarity. [^7] Note that 0 relevant docs samples are usually numeric or N/A types answer entities which results in high accuracy gain.
@@
Human judgement on Model prediction: Incorrect
Insteval judgement on Model prediction: Correct
@@
Human judgement on Model prediction: Correct
Insteval judgement on Model prediction: Incorrect
Table 3: Error instances from Insteval, which do not match human judgement. The examples are taken from NQ and ten samples from each of the five QA models.
Figure 4: Average accuracy against human labels across five QA models, using original answer set and our methods. We separately report the accuracy based on the rarity of the entity, which is measured by the number of its relevant docs in DBpedia (Kandpal et al., 2023).
Environmental footprint
@@ B.4 Case study: The impact of answer set expansion
B.5 Case study: Rationale behind Soft EM
Although our method, which utilizes Soft EM within an expanded answer set, shows reliable performance in two well-known QA datasets–Natural Questions and TriviaQA–it is important to consider the potential for differing results in specialized domains. In well-known datasets, where LLMs like GPT-3.5 might have encountered similar data during training, it is less likely to generate responses with significant contextual errors (e.g. “Joe Biden is not the president of US“), which cannot be deemed correct by matching “Joe Biden”. However, the potential for such errors could increase in specialized domains.
To investigate the effectiveness of Soft EM in these specialized domains, we conduct an additional experiment using the SciQ dataset Welbl et al. (2017), known for its specialized science-based content. We utilized the InstructGPT model (gpt-3.5-turbo-instruct) to answer 100 randomly sampled questions from the SciQ dataset. Each response was manually verified to determine whether the correct entity was identified and whether it was placed within an accurate context.
Out of 100 questions manually analyzed, the QA model generated an incorrect answer containing the correct entity on only 2 questions. These instances were labeled as incorrect by humans despite containing the gold answer because they listed multiple answers, including the correct one. This high level of correlation with human judgment underscores the reliability of Soft EM as an evaluation metric, capable of effectively measuring QA performance across specialized domains.